# Wide-Field, High-Resolution Reconstruction in Computational Multi-Aperture Miniscope Using a Fourier Neural Network

### v1

@@
Wide-Field, High-Resolution Reconstruction in Computational Multi-Aperture Miniscope Using a Fourier Neural Network
[…]
journal: opticajournalarticletype: Research Article{abstract*}
@@ 1 Introduction
Wide-field, high-resolution imaging plays a critical role across diverse scientific disciplines. Yet, traditional microscopy, which relies on single-objective lenses, is encumbered by a fundamental trade-off between field-of-view (FOV) and resolution[1]. To circumvent these constraints, array microscopes have been developed, facilitating rapid, wide-FOV image acquisition[2, 3, 4]. These systems utilize an ensemble of microscopes, each having its own lens and image sensor, to expand the FOV without sacrificing resolution. Despite these advancements, the scalability and miniaturization of array microscopes are limited by their dependence on bulky, specialized image sensor arrays, which pose significant challenges in applications demanding portability and compactness, such as endoscopy[5], on-chip microscopy[6], and in-vivo neural imaging [7].
@@
Departing from conventional array microscopes, our system utilizes a single image sensor to enhance device miniaturization and portability. Unlike other MLA-based multi-aperture imaging designs[11, 12, 13] that employ physical barriers to isolate sub-images captured by individual microlenses, our system opts for an unobstructed arrangement to maximize the effective imaging FOV. This design, however, introduces severe cross-talk between adjacent sub-images on the sensor, i.e. view multiplexing. Furthermore, the single-element MLA generates point spread functions (PSFs) that contain widely separated array foci and exhibit considerable spatial variability, i.e. non-local and spatially-varying PSFs. These complexities pose substantial challenges to achieving high-quality image reconstruction across the extended FOV.
@@
A second challenge in using CNNs for handling non-local, spatially-varying PSFs is the networks’ inherent locality bias [19, 20], a result of the small filter sizes in convolution layers that restrict CNN to capture global information. Although increasing CNN’s depth can broaden the “receptive field”, this approach escalates computational costs and complicates network training. The recently proposed MultiWienerNet [17] addresses these issues by initially applying deconvolution with multiple Wiener filters informed by a set of pre-calibrated, spatially-varying PSFs, followed by refinement through a CNN. Yet, this technique heavily depends on the choice of these initial PSFs that need to be sufficiently representative of the underlying spatially-varying aberrations across the FOV. An alternative approach is the Fourier Neural Network (FourierNet), which overcomes locality bias by applying filters in the Fourier domain [21, 22, 23, 24]. This strategy takes advantage of the global processing capabilities of the Fourier transform, allowing FourierNet to achieve a global receptive field with minimal layers, thus efficiently sidestepping the limitations faced by traditional CNNs. While FourierNet has been successfully applied in spatially invariant imaging systems [22, 23, 24], its ability to address spatially-varying systems remains unexplored.
In this work, we introduce Spatially-Varying FourierNet (SV-FourierNet) to facilitate wide-field, high-resolution reconstruction within our computational multi-aperture miniscope. We show that SV-FourierNet can be applied directly on full-FOV measurements containing 12-million pixels, circumventing the uneven resolution constraints typically associated with patch-based training methods. Furthermore, by employing multiple learnable Fourier filters [Fig. 1(b)], SV-FourierNet is effective in addressing the view multiplexing and highly non-local and spatially-varying array PSFs in our multi-aperture miniscope. The multi-channel Fourier layer in SV-FourierNet effectively performs a blind deconvolution, based on a low-rank imaging model, to process the input measurement. The resulting multi-channel deconvolved outputs are then fused and refined by an enhancement network, yielding the final reconstructed image.
@@
Compared to current existing state-of-the-art deep learning methods[10, 17], SV-FourierNet demonstrates superior reconstruction quality, consistent resolution across the FOV, and enhanced inference speed. Trained entirely on a physics-based simulator, we show experimentally that SV-FourierNet attains a uniform resolution of 7.8 $\mu$m across a 6.5 mm FOV. Additionally, we present wide-field, high-resolution video reconstructions of freely moving C. elegans colonies and imaging of a weakly scattering mouse brain section.
@@ 2 Methods
2.1 Setup of the multi-aperture miniscope
The multi-aperture miniscope is built with off-the-shelf and 3D-printed components. It uses an off-the-shelf plastic 3$\times$3 MLA (1 mm pitch, no. 630, Fresnel Technologies Inc.), forming a finite conjugate imaging geometry with a working distance of $\sim$7.5 mm and an MLA-sensor distance of $\sim$4.5 mm. For green fluorescence imaging, a hybrid filter set is incorporated, combining an interference filter (no. 535/50, Chroma Technology) placed in front of the MLA and a long-pass absorption filter (Wratten color filter no. 12, Edmund Optics) placed after the MLA. It uses a backside-illuminated (BSI) CMOS sensor (IM226, IDS Imaging) with 12 million pixels and 1.85 $\mu$m pixel size, achieving an effective pixel size of 3.7 $\mu$m. See more details in Supplement 1.
@@ 2.2 Spatially-varying imaging model
To expedite the simulation process, sub-images from individual MLs are computed in parallel, allowing the model to generate a measurement within 30 seconds. The sensor noise is sampled from a pre-calibrated mixed Poisson-Gaussian noise model and added as an augmentation to the measurements during the training process, which ensures that the synthetic measurements closely approximate real imaging conditions (See details in Supplement 1).
The spatially-varying imaging model is validated by comparing the simulated and experimental measurements on the same object containing randomly distributed particles, shown in Fig. 2(e) (See details in Supplement 1). The agreement between simulation and experiment, measured by PCC, achieves 0.92, indicating our model’s accuracy. Furthermore, two zoomed-in regions show that the model accurately reproduces the aberrations present in the physical measurement.
To promote generalization of SV-FourierNet, we simulate training data by applying our low-rank model to a diverse set of biological images collected from online repositories[27, 28, 29, 30, 31], including cells, vasculature, and brain sections, as well as synthetically generated fluorescent particles with random size, brightness, and locations, yielding around 5000 training images.
@@ 2.3 SV-FourierNet Design
SV-FourierNet comprises two core modules, including a blind Fourier deconvolution module and an enhancement module, as shown in Fig. 3. The measurement is first Fourier transformed into the frequency domain, followed by element-wise multiplication with multiple learnable complex Fourier filters, whose weights are initialized with a uniform distribution $U$[0, 1]. The filtered results are then inverse Fourier transformed back to the spatial domain, resulting in the initial multi-channel reconstructions. This blind deconvolution approach, in contrast to MultiWienerNet [17], bypasses the need for pre-selecting PSFs and facilitates the efficient learning of a low-rank Fourier basis. Additionally, by leveraging the conjugate symmetry of the Fourier transform, SV-FourierNet halves the number of training parameters required compared to those needed for learning directly in the spatial domain[17]. SV-FourierNet employs nine Fourier filters to collectively account for the spatial variance in our system.
The enhancement module employs a modified Residual Channel Attention Network (RCAN) [32], which includes a down-scaling module, a residual in residual (RIR) block and an up-scaling module. Each RIR block connects two residual groups with a long skip connection, facilitating information fusion across different scales. Channel attention modules embedded within the residual groups dynamically adjust the channel-wise feature weights, essential for synthesizing information from different Fourier channels. To prevent overfitting and to improve generalizability, the enhancement module is designed to be lightweight, with approximately 0.18 million parameters. Additional details about the network implementation is in Supplement 1.
@@
We implement SV-FourierNet in PyTorch on an Nvidia A40 GPU. Training is conducted entirely on a simulated dataset with a batch size of 4. We preprocess the simulated measurement through cropping and padding operations to address the periodic ambiguity characteristic of array PSFs (see details in Supplement 1). Optimization employs the Adam optimizer alongside a cosine annealing scheduler, with the entire training process spanning approximately 48 hours.
@@ 2.4 Computation of the ERF
To assess SV-FourierNet’s ability to encapsulate non-local, spatially-varying aberrations and learn a physically meaningful inverse mapping, we calculate its ERF after the network is trained. This involves gradient backpropagation from a pixel on the reconstructed object to the input measurement[25], effectively measuring the influence of the measurement on a specific object pixel and generating a localized heatmap. In practice, this heatmap depends not only on the network’s trained weights, but also varies with different input objects, making it input-dependent. The ERF is the ensemble average of the heatmaps across a large dataset, effectively making it input-independent. The SV-FourierNet’s ERF is computed using a set of 520 test images.
@@ 3.1 SV-FourierNet’s learned ERF aligns with physical PSFs
SV-FourierNet’s ERFs contain both positive and negative components. The positive ERF components indicate that specific pixels in the measurement lead to an increase in intensity at the reconstructed pixel by SV-FourierNet. Given that PSFs in fluorescence imaging are inherently positive, it is the positive ERF components that parallel the physical PSFs’ function. As shown in Fig. 4, SV-FourierNet’s ERFs are highly localized and closely match the physically measured PSFs across the field. At the FOV center, the ERF aligns with the physical PSF, containing a widely distributed 3 $\times$ 3 foci array, as shown in Fig. 4(a). This demonstrates that SV-FourierNet has learned a global receptive field to synthesize information across the entire measurement and the network’s effective response agrees with the physical response of the imaging system. Near the FOV edges, due to the sensor’s finite size and the system’s limited angular response, the physical PSFs are truncated to 2 $\times$ 3 or 3 $\times$ 2 foci array respectively, as shown in Fig. 4(b) and 4(c). Nevertheless, the ERFs precisely localize the physical foci. This showcases SV-FourierNet’s capability to learn the severe spatial variance in the imaging model, including handling image truncation – a case that often induces numerical artifacts in model-based methods[35]. In addition, when assessing a point outside the SV-FourierNet’s trained FOV, the physical PSF appears as a truncated 2 $\times$ 2 foci array with noticeable contrast loss [Fig. 4(d)]. However, the ERF aligns precisely with this pattern, underscoring SV-FourierNet’s ability to learn a physically accurate inverse mapping. This ensures model generalization even beyond the trained FOV boundaries. Zoomed-in views of the foci regions further elucidate the severe aberrations and spatial variability of the PSFs, along with the precise co-localization between the ERFs and the physical PSFs.
@@
The negative ERF components contribute to SV-FourierNet’s capability to demultiplex overlapping views in the measurement, essential for addressing the view-multiplexing challenge inherent to our system. These components, as detailed in Supplement 1, form a periodic array foci pattern corresponding to the MLA’s periodicity. This pattern arises from the measurement’s ambiguity caused by the MLA’s periodic structure, where displacement by a distance equal to the foci array’s period results in nearly identical measurements. The differences are only due to unique aberrations of the individual microlenses and the image truncation by the system. SV-FourierNet navigates this complexity by leveraging distinct aberration patterns across microlenses, accentuating the accurate object components through positive ERF components while suppressing contributions from ambiguous components via negative ERF components.
3.2 Demonstration of consistent resolution across a wide FOV
@@
To showcase our system’s capability to achieve uniform resolution across a wide FOV, we characterize the resolution at both the FOV center and edges by imaging a fluorescence resolution target, as shown in Fig. 5. We also compare the reconstructions with the raw measurements from the central lens, which underscores the resolution enhancement and FOV expansion enabled by SV-FourierNet. In the central FOV, SV-FourierNet’s reconstruction surpass those from the central lens in both resolution and contrast, illustrating the network’s efficiency in effectively synthesizing information from multiple lenses for high-resolution reconstructions. At the FOV periphery, where the central lens exhibits pronounced aberration and vignetting artifacts, SV-FourierNet robustly reconstructs the target with high resolution, showcasing the network’s capability to account for strong spatially-varying aberrations. The line profiles further affirm SV-FourierNet’s capacity to achieve a consistent 7.8 $\mu$m resolution vertically and horizontally at both central and peripheral FOV regions.
@@
Figure 5: Resolution characterization. SV-FourierNet reconstruction of fluorescent resolution target at the FOV center and edges, compared with measurements from the central lens to highlight the enhanced resolution and extended FOV achieved by our system. Line profiles confirms the system achieves a consistent 7.8 $\mu$m resolution both vertically and horizontally, at both FOV center and edges.

### v2

@@
Abstract
@@ 1 Introduction
Wide-field, high-resolution imaging plays a critical role across diverse scientific disciplines. Yet, traditional microscopy, which relies on single-objective lenses, is encumbered by a fundamental trade-off between field-of-view (FOV) and resolution[1]. To circumvent these constraints, array microscopes have been developed, facilitating rapid, wide-FOV image acquisition[2, 3, 4]. These systems utilize an ensemble of microscopes, each having its own lens and image sensor, to expand the FOV without sacrificing resolution. Despite these advancements, the scalability and miniaturization of array microscopes are limited by their dependence on bulky, specialized image sensor arrays, which pose significant challenges in applications demanding portability and compactness, such as endoscopy[5], on-chip microscopy[6], and in-vivo neural imaging [7]. Recently, lensless imaging techniques have emerged as a compact alternative by employing a phase mask[8, 9] placed directly in front of a CMOS sensor. However, the removal of focusing optics leads to reduced contrast and signal-to-noise ratio (SNR), which restricts their sensitivity for imaging weak fluorescent signals[10, 11].
@@
Departing from conventional array microscopes, our system utilizes a single image sensor to enhance device miniaturization and portability. Unlike other MLA-based multi-aperture imaging designs[15, 16, 17] that employ physical barriers to isolate sub-images captured by individual microlenses, our system opts for an unobstructed arrangement to maximize the effective imaging FOV[13]. This design, however, introduces severe cross-talk between adjacent sub-images on the sensor, i.e. view multiplexing. Furthermore, the single-element MLA generates point spread functions (PSFs) that contain widely separated array foci and exhibit considerable spatial variability, i.e. non-local and spatially-varying PSFs. These complexities pose substantial challenges to achieving high-quality image reconstruction across the extended FOV.
@@
A second challenge in using CNNs for handling non-local, spatially-varying PSFs is the networks’ inherent locality bias [23, 24], a result of the small filter sizes in convolution layers that restrict CNNs from capturing global information. Although increasing CNN’s depth can broaden the “receptive field”, this approach escalates computational costs and complicates network training. The recently proposed MultiWienerNet and MultiFlatNet [21, 22] address these issues by initially applying deconvolution with multiple Wiener filters informed by a set of pre-calibrated, spatially-varying PSFs, followed by refinement through a CNN. Yet, this type of techniques heavily depends on the choice of these initial PSFs that need to be sufficiently representative of the underlying spatially-varying aberrations across the FOV. Our study also shows that this spatial domain learning approach cannot efficiently reconstruct high resolution information due to the implicit bias towards low-frequency features induced by the Wiener filtering process. An alternative approach is the Fourier domain Neural Network (FourierNet), which overcomes locality bias by applying filters directly in the Fourier domain [25, 26, 27, 28]. This strategy explicitly takes advantage of the global processing capabilities of the Fourier transform, allowing FourierNet to achieve a global receptive field with minimal layers, thus efficiently sidestepping the limitations faced by traditional CNNs. While FourierNet has been successfully applied in spatially invariant imaging systems [26, 27, 28], its ability to address spatially-varying systems remains unexplored.
In this work, we introduce Spatially-Varying FourierNet (SV-FourierNet) to facilitate wide-field, high-resolution reconstruction within our computational multi-aperture miniscope. We show that SV-FourierNet can be applied directly on full-FOV measurements containing 12-million pixels, circumventing the uneven resolution constraints typically associated with patch-based training methods. Furthermore, by employing multiple learnable Fourier filters [Fig. 1(b)], SV-FourierNet is effective in addressing the view multiplexing and highly non-local and spatially-varying array PSFs in our multi-aperture miniscope. The multi-channel Fourier layer in SV-FourierNet effectively performs a blind deconvolution, based on a low-rank imaging model, to process the input measurement. The resulting multi-channel deconvolved outputs are then fused and refined by an enhancement network, yielding the final reconstructed image. Our study shows that this Fourier domain learning approach can more efficiently capture high resolution information as compared to the spatial domain learning methods.
@@
Compared to current existing state-of-the-art model-based and deep-learning-based methods[14, 21, 22], SV-FourierNet demonstrates superior reconstruction quality, consistently higher resolution across the FOV, and enhanced inference speed. Trained entirely on a physics-based simulator, we show experimentally that SV-FourierNet attains a uniform resolution of 7.8 $\mu$m across a 6.5 mm FOV. The network also provides consistent reconstruction quality across an extended depth-of-field (EDOF) over a 100 $\mu$m depth range. Additionally, we present wide-field, high-resolution video reconstructions of freely moving C. elegans colonies and imaging of a weakly scattering mouse brain section.
@@ 2 Methods
2.1 The multi-aperture miniscope
The multi-aperture miniscope is built with off-the-shelf and 3D-printed components. It uses an off-the-shelf plastic 3$\times$3 MLA (1 mm pitch, no. 630, Fresnel Technologies Inc.), forming a finite conjugate imaging geometry with a working distance of $\sim$7.5 mm and an MLA-sensor distance of $\sim$4.5 mm. Each lens within the MLA captures an image of a distinct and limited sub-FOV, while the whole array covers an extended FOV (see more details in Section 1 in Supplement 1). For green fluorescence imaging, a hybrid filter set is incorporated, combining an interference filter (no. 535/50, Chroma Technology) placed in front of the MLA and a long-pass absorption filter (Wratten color filter no. 12, Edmund Optics) placed after the MLA. It uses a backside-illuminated (BSI) CMOS sensor (IM226, IDS Imaging) with 12 million pixels and 1.85 $\mu$m pixel size, achieving an effective pixel size of 2.8 $\mu$m. To validate the reconstruction results from our system, we set up a widefield microscope to record fluorescence simultaneously with our miniscope. See more details in Section 9A-B in Supplement 1.
@@ 2.2 Spatially-varying imaging model
To expedite the simulation process, sub-images from individual MLs are computed in parallel, allowing the model to generate a measurement within 30 seconds. The sensor noise is sampled from a pre-calibrated mixed Poisson-Gaussian noise model. The noise is added to the measurements in the data loader as a form of “online” data augmentation during the training process, which ensures that the synthetic measurements closely approximate real imaging conditions (see details in Section 3B in Supplement 1).
The spatially-varying imaging model is validated by comparing the simulated and experimental measurements on the same object containing randomly distributed particles, shown in Fig. 2(e) (See details in Section 3A in Supplement 1). The agreement between simulation and experiment, measured by PCC, achieves 0.92, indicating our model’s accuracy. Furthermore, two zoomed-in regions show that the model accurately reproduces the aberrations present in the physical measurement.
To promote generalization of SV-FourierNet, we simulate training data by applying our low-rank model to a diverse set of biological images collected from online repositories[31, 32, 33, 34, 35], including cells, vasculature, and brain sections, as well as synthetically generated fluorescent particles with random size, brightness, and locations, yielding around 5000 training images (see details in Section 3C in Supplement 1).
@@ 2.3 SV-FourierNet Design
SV-FourierNet comprises two core modules, including a blind Fourier deconvolution module and an enhancement module, as shown in Fig. 3. The measurement is first Fourier transformed into the frequency domain, followed by element-wise multiplication with multiple learnable complex Fourier filters, whose weights are initialized with a uniform distribution $U$[0, 1]. The filtered results are then inverse Fourier transformed back to the spatial domain, resulting in the initial multi-channel reconstructions.
The Fourier-domain blind deconvolution approach in SV-FourierNet, in contrast to MultiWienerNet / MultiFlatNet [21, 22], bypasses the need for pre-selecting PSFs and facilitates the efficient learning of a low-rank Fourier basis. Additionally, by leveraging the conjugate symmetry of the Fourier transform, SV-FourierNet halves the number of training parameters required compared to those needed for learning directly in the spatial domain[21, 22]. SV-FourierNet employs nine Fourier filters to collectively account for the spatial variance in our system. The number of filters is crucial as it dictates the network’s capacity for nonlinearity. We have conducted an ablation study on the number of Fourier filters, which revealed that increasing the number of filters improves resolution and reconstruction quality but at a cost of memory consumption. Given the trade-off between network performance and computational cost, we have opted for nine channels to facilitate the spatially varying reconstruction (see details in Section 4C in Supplement 1).
The enhancement module employs a modified Residual Channel Attention Network (RCAN) [36], which includes a down-scaling module, a residual in residual (RIR) block and an up-scaling module. Each RIR block connects two residual groups with a long skip connection, facilitating information fusion across different scales. Channel attention modules embedded within the residual groups dynamically adjust the channel-wise feature weights, essential for synthesizing information from different Fourier channels. To prevent overfitting and to improve generalizability, the enhancement module is designed to be lightweight, with approximately 0.18 million parameters. Additional details about the network implementation is in Section 4B in Supplement 1. In addition, cropping and padding operations are implemented in the data loader to address the periodic ambiguity characteristic of array PSFs (see details in Section 4A in Supplement 1).
@@
We implement SV-FourierNet in PyTorch on an Nvidia A40 GPU. Training is conducted entirely on a simulated dataset with a batch size of 4. Optimization employs the Adam optimizer alongside a cosine annealing scheduler, with the entire training process spanning approximately 48 hours (See detailed loss curve in Section 4D in Supplement 1).
@@ 2.4 Computation of the ERF
To assess SV-FourierNet’s ability to encapsulate non-local, spatially-varying aberrations and learn a physically meaningful inverse mapping, we calculate its ERF after the network is trained. This involves gradient backpropagation from a pixel on the reconstructed object to the input measurement[29], effectively measuring the influence of the measurement on a specific object pixel and generating a localized heatmap. In practice, this heatmap depends not only on the network’s trained weights, but also varies with different input objects, making it input-dependent. The ERF is the ensemble average of the heatmaps across a large dataset, effectively making it input-independent. The SV-FourierNet’s ERF is computed using a set of 520 test images (See detailed implementation in Section 5A in Supplement 1).
@@
2.5 Model-based reconstruction algorithms
We derive and implement an LSV model-based algorithm based on solving a regularized least-squares problem using the alternating direction method of multiplier (ADMM) algorithm, as detailed in Section 2 in Supplement 1.
@@ 3.1 SV-FourierNet’s learned ERF aligns with physical PSFs
SV-FourierNet’s ERFs contain both positive and negative components. The positive ERF components indicate that specific pixels in the measurement lead to an increase in intensity at the reconstructed pixel by SV-FourierNet. Given that PSFs in fluorescence imaging are inherently positive, it is the positive ERF components that parallel the physical PSFs’ function. As shown in Fig. 4, SV-FourierNet’s ERFs are highly localized and closely match the physically measured PSFs across the field. At the FOV center, the ERF aligns with the physical PSF, containing a widely distributed 3 $\times$ 3 foci array, as shown in Fig. 4(a). This demonstrates that SV-FourierNet has learned a global receptive field to synthesize information across the entire measurement and the network’s effective response agrees with the physical response of the imaging system. Near the FOV edges, due to the sensor’s finite size and the system’s limited angular response, the physical PSFs are truncated to 2 $\times$ 3 or 3 $\times$ 2 foci array respectively, as shown in Fig. 4(b) and 4(c). Nevertheless, the ERFs precisely localize the physical foci. This showcases SV-FourierNet’s capability to learn the severe spatial variance in the imaging model, including handling image truncation – a case that often induces numerical artifacts in model-based methods[39]. In addition, when assessing a point outside the SV-FourierNet’s trained FOV, the physical PSF appears as a truncated 2 $\times$ 2 foci array with noticeable contrast loss [Fig. 4(d)]. However, the ERF aligns precisely with this pattern, underscoring SV-FourierNet’s ability to learn a physically accurate inverse mapping. This ensures model generalization even beyond the trained FOV boundaries. Zoomed-in views of the foci regions further elucidate the severe aberrations and spatial variability of the PSFs, along with the precise co-localization between the ERFs and the physical PSFs. A detailed quantitative analysis and the evolution of ERF throughout the training process are provided in Section 5B-C in Supplement 1.
@@
The negative ERF components demonstrate SV-FourierNet’s capability to suppress aberrations and demultiplex overlapping views in the measurement. These negative components, as detailed in Section 5D in Supplement 1, either surround the positive ERF to enhance resolution or create an additional periodic array foci pattern that corresponds to the MLA’s periodicity. This pattern arises from the measurement’s ambiguity caused by the MLA’s periodic structure, where displacement by a distance equal to the foci array’s period results in nearly identical measurements. The differences are only due to unique aberrations of the individual microlenses and the image truncation by the system. SV-FourierNet navigates this complexity by leveraging distinct aberration patterns across microlenses, accentuating the accurate object components through positive ERF components while suppressing contributions from ambiguous components via negative ERF components.
Finally, to elucidate the physical mapping function learned by the blind deconvolution module of SV-FourierNet, excluding the nonlinearity induced by the enhancement network module, we compute and visualize the equivalent spatial domain filters by transforming the learned Fourier filters to the real space, as detailed in Section 6A of Supplement 1. Notably, the visualization reveals that almost all the spatial domain filters exhibit 3 $\times$ 5 foci, which contrasts with the physical PSFs that contain 3 $\times$ 3 (or fewer) foci. To further understand this result, we compute the pseudo-inverse of the first nine basis PSFs derived via TSVD from experimentally calibrated PSFs, as detailed in Section 6B of Supplement 1. These basis filters from our physical model show a high degree of similarity to the filters learned by SV-FourierNet, each also displaying a 3 $\times$ 5 array of foci. Thus, our physical interpretation of the learned Fourier filters is a low-rank approximation of the pseudo-inverse of the LSV imaging model.
3.2 Demonstration of consistent resolution across a wide FOV and an EDOF
@@
To showcase our system’s capability to achieve uniform resolution across a wide FOV, we characterize the resolution at both the FOV center and edges by imaging a fluorescence resolution target, as shown in Fig. 5(a). We also compare the reconstructions with the raw measurements from the central lens, which underscores the resolution enhancement and FOV expansion enabled by SV-FourierNet. In the central FOV, SV-FourierNet’s reconstruction surpass those from the central lens in both resolution and contrast, illustrating the network’s efficiency in effectively synthesizing information from multiple lenses for high-resolution reconstructions. At the FOV periphery, where the central lens exhibits pronounced aberration and vignetting artifacts, SV-FourierNet robustly reconstructs the target with high resolution, showcasing the network’s capability to account for strong spatially-varying aberrations. The line profiles further affirm SV-FourierNet’s capacity to achieve a consistent 7.8 $\mu$m resolution vertically and horizontally at both central and peripheral FOV regions (See resolution characterization at additional FOV positions in Section 8C in Supplement 1).
In addition, we characterize the reconstructed resolution of SV-FourierNet by computing the power spectral density (PSD). This analysis is performed on patches at different radial displacements and across various experimental samples. The PSD exhibits minimal variation, indicating that SV-FourierNet achieves consistent resolution across the FOV and demonstrates robustness to different experimental samples (see details in Section 8D of Supplement 1).
[…]
@@
Figure 5: Resolution characterization. (a) SV-FourierNet reconstruction of fluorescent resolution target at the FOV center and edges, compared with measurements from the central lens to highlight the enhanced resolution and extended FOV achieved by our system. Line profiles confirms the system achieves a consistent 7.8 $\mu$m resolution both vertically and horizontally, at both FOV center and edges. (b) SV-FourierNet maintains a consistent resolution within a 100 $\mu$m depth range. Beyond this range, resolution begins to degrade and background artifacts appear, as indicated by the red arrow.